# Colegiado de Física, Departamento de Ciências Exatas e da Terra, Universidade do Estado da Bahia Rua Silveira Martins, Cabula, 41150000, Salvador, Bahia, Brazil

### 24 Mar 2026 v1

@@
Inequalities for the Tsallis q-entropy and Information Theory 111matrindade@uneb.br Marco A. S. Trindade,
@@
In this work, we derive information-theoretic properties for a modified Tsallis entropy, hereinafter referred to as $q$-entropy. We introduce the notions of joint $q$-entropy, conditional $q$-entropy, relative $q$-entropy, conditional mutual $q$-information, and establish several inequalities analogous to those of classical information theory. Within the context of Markov chains, these results are employed to prove a version of the second law of thermodynamics. Furthermore, we investigate the maximum entropy method in this setting. Finally, we prove a Tsallis version of the Shannon-McMillan-Breiman theorem.
@@ 1 Introduction
Over the past decades, Tsallis statistics has been successfully applied to a broad class of physical systems in which long-range interactions, long-term memory effects, or multifractal phase-space structures play a relevant role [7, 8]. Such systems frequently arise in statistical physics, plasma physics, turbulence, gravitational systems, and other complex systems [9, 10, 11], where the assumptions underlying Boltzmann-Gibbs statistics are not fully satisfied [12]. In this scenario, the nonadditive nature of Tsallis entropy provides a theoretical framework for describing anomalous scaling behaviors and generalized thermodynamic relations, which are not adequately captured by conventional entropic measures [12].
@@
The paper is organized as follows. Section 2 contains the basic definitions and inequalities. In Section 3 we perform a stochastic analysis and we derive a second law of thermodynamics in the context of Markov chain. In Section 4 we apply the maximum entropy method in this scenario. Section 6 is devoted to presenting a Tsallis version of the Shannon-McMillan-Breiman theorem. Section 7 is dedicated to conclusions and perspectives. Basic probabilistic concepts relevant to this study are summarized in Appendix A.
@@ 2 Basic Results
Proof. Fix $q\neq 1$ and let $t\in(0,1]$. By definition,
$\ln_{q}t=\frac{t^{1-q}-1}{1-q}.$
@@
$\ln_{q}t\leq 0,$
with equality if and only if $t=1$.
Since $p(x)\geq 0$ for all $x\in\chi$, it follows that
$p(x)\ln_{q}p(x)\geq 0,$
@@ 4 Maximum entropy method for the q-entropy
The Maximum Entropy method was originally formulated by Edwin T. Jaynes [24, 2]. Its central idea is to select, among all probability distributions compatible with a given set of macroscopic constraints (such as known expectation values). In the classical framework, the Shannon entropy is employed, and its maximization naturally leads to the familiar exponential distributions of the statistical mechanics developed by Boltzmann and Gibbs [24]. In this context, we can apply the maximum entropy method [24, 2] for $q$- entropy with the constraints:
@@
$H-H^{\prime}$ $=$ $-\sum_{i=1}^{n}[1+(1-q)\ln_{q}p_{i}]f_{i}\ln_{q}\left(\frac{f_{i}}{p_{i}}\right)$ (48)
@@
In this section, we present a version of Shannon-McMillan-Breiman theorem [2] within the nonadditive entropic scenario considered in this work. The result characterizes the asymptotic behavior of information associated with stationary stochastic processes and establishes a qq–version of the asymptotic equipartition property for general ergodic process. In this section, we will consider $1/2<q<1$. Some basic probabilistic concepts relevant to this section are summarized in Appendix A
@@
Let $\{X_{n}\}_{n\in\mathbb{Z}}$ be a stationary ergodic stochastic process over a finite alphabet $\mathcal{X}$. Fix $q<1$ and define the $q$-logarithm:
$\ln_{q}(x)=\frac{x^{1-q}-1}{1-q}.$ (49)
Define the non-extensive conditional entropy rate:
@@
$\limsup_{n\to\infty}\Big(-\frac{1}{n}\ln_{q}p(X_{0}^{\,n-1}|X_{-1}^{\infty})\Big)\leq H_{q,\infty}.$ (53)
@@
$-\frac{1}{n}\ln_{q}p(X_{0}^{n-1}|X_{-1}^{\infty})\leq\frac{1}{n}\sum_{i=0}^{n-1}Z_{i}.$ (58)
@@
$\limsup_{n\to\infty}\Big(-\frac{1}{n}\ln_{q}p(X_{0}^{n-1}|X_{-1}^{\infty})\Big)\leq H_{q,\infty}.$ (59)
@@
$\limsup_{n\to\infty}\Big(-\frac{1}{n}\ln_{q}p(X_{0}^{n-1}|X_{-1}^{\infty})\Big)\leq-\sum_{x_{-\infty}^{-1}}p(x_{-\infty}^{-1})\sum_{x_{0}\in\mathcal{X}}p(x_{0}\mid x_{-\infty}^{-1})\frac{p(x_{0}\mid x_{-\infty}^{-1})^{1-q}-1}{1-q}.$ (60)
@@
Let $\{X_{n}\}_{n\in\mathbb{Z}}$ be a stationary stochastic process with finite alphabet. Assume in addition that there exists a constant $C>0$ such that for all blocks $x_{0}^{n-1}$,
$p(x_{0}^{n-1})\leq C\,p_{k}(x_{0}^{n-1}).$ (62)
@@
Proof. Let $A$ be the support of $p(X_{0}^{n-1})$. Then
$\mathbb{E}\left[\frac{p(X_{0}^{n-1}\mid X_{-\infty}^{-1})}{p(X_{0}^{n-1})}\right]$ $=$ $\mathbb{E}\left[\mathbb{E}\left(\frac{p(X_{0}^{n-1}\mid X_{-\infty}^{-1})}{p(X_{0}^{n-1})}\Big|X_{-\infty}^{-1}\right)\right]$ (65)
$=$ $\mathbb{E}\left[\sum_{x_{0}^{n-1}\in A}p(x_{0}^{n-1}\mid X_{-\infty}^{-1})\right]$ (66)
$\leq$ $1.$ (67)
Similarly, using the additional hypothesis, we have
$\mathbb{E}\left[\frac{p(X_{0}^{n-1})}{p_{k}(X_{0}^{n-1})}\right]$ $=$ $\sum_{x_{0}^{n-1}\in A}p(x_{0}^{n-1})\frac{p(x_{0}^{n-1})}{p_{k}(x_{0}^{n-1})}$ (68)
$\leq$ $\sum_{x_{0}^{n-1}\in A}p(x_{0}^{n-1})C$ (69)
$=$ $C.$ (70)
@@
$\Pr\left\{\frac{p(X_{0}^{n-1}\mid X_{-\infty}^{-1})}{p(X_{0}^{n-1})}\geq t_{n}\right\}$ $\leq$ $\frac{1}{t_{n}},$ (71)
$\Pr\left\{\frac{p(X_{0}^{n-1})}{p_{k}(X_{0}^{n-1})}\geq t_{n}\right\}$ $\leq$ $\frac{C}{t_{n}}.$ (72)
@@
$\Pr\left\{\frac{1}{n}\ln_{q}\frac{p(X_{0}^{n-1}\mid X_{-\infty}^{-1})}{p(X_{0}^{n-1})}\geq\frac{1}{n}\ln_{q}t_{n}\right\}$ $\leq$ $\frac{1}{t_{n}},$ (73)
$\Pr\left\{\frac{1}{n}\ln_{q}\frac{p(X_{0}^{n-1})}{p_{k}(X_{0}^{n-1})}\geq\frac{1}{n}\ln_{q}t_{n}\right\}$ $\leq$ $\frac{C}{t_{n}}.$ (74)
@@
Let $\{X_{n}\}_{n\in\mathbb{Z}}$ be a stationary ergodic stochastic process over a finite alphabet $\mathcal{X}$. Assume in addition that there exists a constant $C>0$ such that for all blocks $x_{0}^{n-1}$,
$p(x_{0}^{n-1})\leq C\,p_{k}(x_{0}^{n-1}).$ (124)
and
@@
$\liminf_{n\to\infty}-\frac{1}{n}\ln_{q}p(X_{0}^{n-1})\geq\liminf_{n\to\infty}-\frac{1}{n}\ln_{q}p_{k}(X_{0}^{n-1})=H_{k}$ (130)
@@
Although the nonadditive parameter $q$ gives rise to further technical difficulties, the result holds rigorously for $1/2<q<1$ provided that condition (124) is satisfied.
@@ 6 Conclusions
In this work, we have introduced definitions and derived information-theoretic inequalities based on a modified Tsallis entropy, which we argue provides a more natural generalization of Shannon entropy. We defined the notions of joint $q$-entropy, conditional $q$-entropy, relative $q$-entropy, and conditional mutual $q$-information, following an approach different from that of Ref. [16], and established several inequalities analogous to those of classical information theory. These results are shown to hold, in general, for $0<q<1$. This restriction may be relaxed, leading to a distinct set of inequalities that no longer preserve a direct analogy with Shannon’s context. Within this scenario, the information-theoretic results developed here were employed, in the context of Markov chains, to prove a version of the second law of thermodynamics. We also applied the formalism to the maximum entropy method. In addition, we proved a Tsallis version of the Shannon-McMillan-Breiman theorem for $1/2<q<1$. As perspectives for future work, potential applications in the context of fractals and multifractals may be explored through the information dimension $lim_{\epsilon\rightarrow 0}\frac{-\langle ln_{q}p\rangle}{ln_{q}1/\epsilon}$, where $\epsilon$ stands for the scaling factor.
@@ Appendix A Probabilistic Tools for Finite Alphabet Sources
In this appendix, we summarize fundamental probabilistic concepts that are frequently used in information theory for processes over finite alphabets. We focus on ergodic transformations, the ergodic theorem for stationary sources, and almost sure convergence results such as the Borel-Cantelli lemma and the convergence dominated theorem. These tools provide the theoretical foundation for results like the asymptotic equipartition property and the Shannon-McMillan-Breiman theorem. In the following, we closely follow the presentations in References [2, 25, 26].
A.1 Ergodic Transformations
@@
Such a transformation is called measure-preserving. A measure-preserving transformation $T$ is ergodic if every $T$-invariant event AA (i.e., $TA=A$) satisfies
@@
A.2 Ergodic Theorem for Stationary Sources
@@ A.3 Borel-Cantelli Lemma
Lemma 7 (Borel-Cantelli)

### 28 Mar 2026 v3

@@
Information-theoretic structure for the Tsallis q-entropy in statistical physics 111matrindade@uneb.br Marco A. S. Trindade,
@@
In this work, we derive information-theoretic properties for a modified Tsallis entropy, hereinafter referred to as $q$-entropy. We introduce the notions of joint $q$-entropy, conditional $q$-entropy, relative $q$-entropy, conditional mutual $q$-information, and establish several inequalities analogous to those of classical information theory. Within the context of Markov chains, these results are employed to prove a version of the second law of thermodynamics. Furthermore, we investigate the maximum entropy method in this setting. Finally, we prove a Tsallis version of the Shannon-McMillan-Breiman theorem and discuss the implications of these results in nonextensive statistical physics.
@@ 1 Introduction
Over the past decades, Tsallis statistics has been successfully applied to a broad class of physical systems in which long-range interactions, long-term memory effects, or multifractal phase-space structures play a relevant role [7, 8]. Such systems frequently arise in statistical physics, plasma physics, turbulence, gravitational systems, and other complex systems [9, 10, 11, 12], where the assumptions underlying Boltzmann-Gibbs statistics are not fully satisfied [13]. In this scenario, the nonadditive nature of Tsallis entropy provides a theoretical scheme for describing anomalous scaling behaviors and generalized thermodynamic relations, which are not adequately captured by conventional entropic measures [13].
@@
The paper is organized as follows. Section 2 contains the basic definitions and inequalities. In Section 3 we perform a stochastic analysis and we derive a second law of thermodynamics in the context of Markov chain. In Section 4 we apply the maximum entropy method in this scenario. Section 6 is devoted to presenting a Tsallis version of the Shannon-McMillan-Breiman theorem. Section 7 is dedicated to conclusions and perspectives. Basic probabilistic concepts relevant to this study are present in Appendix A.
@@ 2 Basic Results
Proof. By definition, with $t\in(0,1]$, we have
$\ln_{q}t=\frac{t^{1-q}-1}{1-q}.$ (5)
@@
$\ln_{q}t\leq 0,$ (6)
with equality if and only if $t=1$. Since $p(x)\geq 0$ for all $x\in\chi$, it follows that
$-\,p(x)\ln_{q}p(x)\geq 0,$ (7)
@@ 4 Maximum entropy method for the q-entropy
The Maximum Entropy method was originally formulated by Edwin T. Jaynes [25, 2]. Its central idea is to select, among all probability distributions compatible with a given set of macroscopic constraints (such as known expectation values). In the classical scenario, the Shannon entropy is employed, and its maximization naturally leads to the familiar exponential distributions of the statistical mechanics developed by Boltzmann and Gibbs [25]. In this context, we can apply the maximum entropy method [25, 2] for $q$- entropy with the constraints:
@@
$H_{q}-H_{q}^{\prime}$ $=$ $-\sum_{i=1}^{n}[1+(1-q)\ln_{q}p_{i}]f_{i}\ln_{q}\left(\frac{f_{i}}{p_{i}}\right)$ (54)
@@
In this section, we present a version of Shannon-McMillan-Breiman theorem [2] within the nonadditive entropic scenario considered in this work. The result characterizes the asymptotic behavior of information associated with stationary stochastic processes and establishes a qq–version of the asymptotic equipartition property for general ergodic process. In this section, we consider $1/2<q<1$, corresponding to systems in which rare microstates are moderately amplified but do not dominate, consistent with ergodic behavior and effective additivity at large scales. Some basic probabilistic concepts relevant to this section are present in Appendix A
@@
Let $\{X_{n}\}_{n\in\mathbb{Z}}$ be a stationary ergodic stochastic process over a finite alphabet $\mathcal{X}$ and $1/2q<1$. Define the non-extensive conditional entropy rate:
@@
$\limsup_{n\to\infty}\Big(-\frac{1}{n}\ln_{q}p(X_{0}^{\,n-1}|X_{-\infty}^{-1})\Big)\leq H_{q,\infty}.$ (58)
@@
$-\frac{1}{n}\ln_{q}p(X_{0}^{n-1}|X_{-\infty}^{-1})\leq\frac{1}{n}\sum_{i=0}^{n-1}Z_{i}.$ (63)
@@
$\limsup_{n\to\infty}\Big(-\frac{1}{n}\ln_{q}p(X_{0}^{n-1}|X_{-\infty}^{-1})\Big)\leq H_{q,\infty}.$ (64)
@@
$\limsup_{n\to\infty}\Big(-\frac{1}{n}\ln_{q}p(X_{0}^{n-1}|X_{-\infty}^{-1})\Big)\leq-\sum_{x_{-\infty}^{-1}}p(x_{-\infty}^{-1})\sum_{x_{0}\in\mathcal{X}}p(x_{0}\mid x_{-\infty}^{-1})\frac{p(x_{0}\mid x_{-\infty}^{-1})^{1-q}-1}{1-q}.$ (65)
@@
Let $\{X_{n}\}_{n\in\mathbb{Z}}$ be a stationary stochastic process with finite alphabet. Assume in addition that there exists constants $C>0$ and $C^{\prime}>0$ such that
$\sup_{n\geq 1}\mathbb{E}\left[\frac{p(X_{0}^{n-1}\mid X_{-\infty}^{-1})}{p(X_{0}^{n-1})}\right]<C$ (67)
and
$\sup_{n\geq 1}\mathbb{E}\left[\frac{p(X_{0}^{n-1})}{p_{k}(X_{0}^{n-1})}\right]<C^{\prime}$ (68)
@@
Proof.
@@
$\Pr\left\{\frac{p(X_{0}^{n-1}\mid X_{-\infty}^{-1})}{p(X_{0}^{n-1})}\geq t_{n}\right\}$ $\leq$ $\frac{C}{t_{n}},$ (71)
$\Pr\left\{\frac{p(X_{0}^{n-1})}{p_{k}(X_{0}^{n-1})}\geq t_{n}\right\}$ $\leq$ $\frac{C^{\prime}}{t_{n}}.$ (72)
@@
$\Pr\left\{\frac{1}{n}\ln_{q}\frac{p(X_{0}^{n-1}\mid X_{-\infty}^{-1})}{p(X_{0}^{n-1})}\geq\frac{1}{n}\ln_{q}t_{n}\right\}$ $\leq$ $\frac{C}{t_{n}},$ (73)
$\Pr\left\{\frac{1}{n}\ln_{q}\frac{p(X_{0}^{n-1})}{p_{k}(X_{0}^{n-1})}\geq\frac{1}{n}\ln_{q}t_{n}\right\}$ $\leq$ $\frac{C^{\prime}}{t_{n}}.$ (74)
@@
The supremum conditions (67) and (68) provide uniform control over block probability ratios and may be satisfied in stationary processes with memory decay, allowing the asymptotic analysis.
@@
The assumption $T_{3}$ ensures that the higher-order interaction terms arising from the nonadditivity of the $q$-logarithm remain negligible at the macroscopic scale. This condition can be interpreted as a weak-interaction requirement and is satisfied in scenarios where correlations are sufficiently controlled.
@@
Let $\{X_{n}\}_{n\in\mathbb{Z}}$ be a stationary ergodic stochastic process over a finite alphabet $\mathcal{X}$. Assume in addition that
@@
$\liminf_{n\to\infty}-\frac{1}{n}\ln_{q}p(X_{0}^{n-1})\geq\liminf_{n\to\infty}-\frac{1}{n}\ln_{q}p_{k}(X_{0}^{n-1})\geq H_{q,k}$ (129)
@@
The conditions (124) and (125) impose a consistency requirement between infinite and finite conditioning and ensure a regular behavior of block probabilities.
Although the nonadditive parameter $q$ introduces additional technical difficulties, the result remains rigorously valid for $1/2<q<1$, provided that the assumptions of Theorem 6 are satisfied. As discussed in [28], the typical set can be identified with the set of accessible microstates, leading to an asymptotic equiprobability of these states. In the present work, this result is extended to the nonextensive framework, yielding a Tsallis-type generalization in which this equiprobability holds in a $q$-deformed context. The assumptions of Theorem 6 ensure this behavior and can be verified in concrete classes of stochastic processes, providing a suitable formulation for the analysis.
@@ 6 Conclusions
In this work, we have introduced definitions and derived information-theoretic inequalities based on a modified Tsallis entropy, which we argue provides a more natural generalization of Shannon entropy. We defined the notions of joint $q$-entropy, conditional $q$-entropy, relative $q$-entropy, and conditional mutual $q$-information, following an approach different from that of Ref. [17], and established several inequalities analogous to those of classical information theory. These results are shown to hold, in general, for $0<q<1$. This restriction may be relaxed, leading to a distinct set of inequalities that no longer preserve a direct analogy with Shannon’s context. Within this scenario, the information-theoretic results developed here were employed, in the context of Markov chains, to prove a version of the second law of thermodynamics. We also applied the formalism to the maximum entropy method. In addition, we proved a Tsallis version of the Shannon-McMillan-Breiman theorem for $1/2<q<1$ and discuss its implications in nonextensive statistical physics, in particular its application to the asymptotic equiprobability of accessible microstates. As perspectives for future work, potential applications in the context of fractals and multifractals may be explored through the information dimension $lim_{\epsilon\rightarrow 0}\frac{-\langle ln_{q}p\rangle}{ln_{q}1/\epsilon}$, where $\epsilon$ stands for the scaling factor.
@@ Appendix A Probabilistic tools for finite alphabet sources
In this appendix, we present fundamental probabilistic concepts that are frequently used in information theory for processes over finite alphabets. We focus on ergodic transformations, the ergodic theorem for stationary sources, and almost sure convergence results such as the Borel-Cantelli lemma and the convergence dominated theorem. These tools provide the theoretical foundation for results like the asymptotic equipartition property and the Shannon-McMillan-Breiman theorem. In the following, we closely follow the References [2, 26, 27].
@@
Such a transformation is called measure-preserving. A measure-preserving transformation $T$ is ergodic if every $T$-invariant event $A$ (i.e., $TA=A$) satisfies
@@
Theorem 7 (Birkhoff ergodic theorem)
@@
Lemma 7 (Borel-Cantelli lemma)